Helping my kids with homework and need to refresh my memory on this stuff!
Solve x^2-12x=0

Answer by Alan3354(69443)


Solve x^2-12x=0
--------------
x*(x-12) = 0
Either x = 0, or x-12 = 0
---
--> x = 0, x = 12